# Energy-Efficient Transformer Inference: Optimization Strategies for Time Series Classification

### v3

@@ Abstract
The increasing computational demands of transformer models in time series classification necessitate effective optimization strategies for energy-efficient deployment. Our study presents a systematic investigation of optimization techniques, focusing on structured pruning and quantization methods for transformer architectures. Through extensive experimentation on three distinct datasets (RefrigerationDevices, ElectricDevices, and PLAID), we quantitatively evaluate model performance and energy efficiency across different transformer configurations. Our experimental results demonstrate that static quantization reduces energy consumption by 29.14% while maintaining classification performance, and L1 pruning achieves a 63% improvement in inference speed with minimal accuracy degradation. Our findings provide valuable insights into the effectiveness of optimization strategies for transformer-based time series classification, establishing a foundation for efficient model deployment in resource-constrained environments.
Introduction
@@
Recent advancements in transformer architectures have significantly advanced time series analysis across various domains, including healthcare, finance, and predictive maintenance, enabling more accurate forecasting, anomaly detection, and decision-making processes [9]. The ability of AI models to process sequential data efficiently has enabled substantial improvements in these fields, allowing for more accurate forecasting, anomaly detection, and decision-making processes [25]. At the core of these developments are transformer-based architectures, which have demonstrated superior performance over traditional models, such as recurrent neural networks (RNNs) and long short-term memory (LSTM) networks, by leveraging self-attention mechanisms to capture long-range dependencies in temporal data [31]. Despite their effectiveness, transformers are computationally expensive, making them less viable for real-time and edge-based applications due to their high energy consumption and memory footprint. Furthermore, the growing carbon footprint associated with transformer training and inference has raised significant concerns regarding sustainability and deployment feasibility in resource-constrained environments [22]. With the exponentially increasing usage of AI, the carbon footprint of massive models has been a topic of increasing worry. At the same time, deep learning model’s runaway scaling, such as huge transformers and diffusion models, has induced unmatched computation costs, which are enormous and require a lot of power to operate and train, directly contributing to the increase in global carbon emissions. Consequently, feeding environmental concerns such as global warming. If not checked, the long-term growth of AI systems will hugely frustrate efforts at sustainability and drain energy resources.
@@
focusing to encompass computation overheads, researchers are developing strategies not to compromise forecasting performance. Enhancing model efficiency, in addition to techniques like knowledge distillation, pruning, quantization, and even hardware-based architectures with increased energy efficiency, should be a pursuit by research professionals. Furthermore, green AI technologies, i.e., renewable energy for data centers and minimizing the algorithms to consume less energy, need to be transitioned in order to make AI a force for good and not an environmental issue.
@@
Transformers in Time Series Forecasting and Classification
@@
Model Compression Strategies
@@
Optimizing Transformer Inference
@@
Transformer Model Implementation
Our study incorporates the Vision Transformer (ViT) model [5], which has been increasingly adapted for time series tasks due to its ability to capture long-range dependencies efficiently [13]. Our model architecture comprises several key components. First, the Time Series Patch Embedding Layer transforms raw time series data into fixed-size patches using a one-dimensional convolutional layer, with positional encoding added to maintain the sequential order. Next, the multi-head self-attention mechanism allows the model to focus on different segments of the time series simultaneously, enhancing its ability to learn complex relationships. Each encoder block contains position-wise feed-forward layers with ReLU activations and dropout layers to reduce overfitting. To ensure stable training and mitigate the risk of vanishing gradients, the architecture incorporates layer normalization and residual connections. Finally, a fully connected softmax output layer classifies each sequence into its corresponding device category.
@@ 3.4.2 Quantization
where $Q$ denotes the quantization factor. Reducing numerical precision can lead to substantial power savings while maintaining a minimal impact on model performance. In our study, we employ two distinct quantization strategies to decrease memory footprint and computational overhead while maintaining high model accuracy.
First, we apply Post-Training Quantization (PTQ). It reduces the precision of model weights and activations after training by converting 32-bit floating-point numbers into lower-bit representations. This process significantly decreases the model’s memory requirements and computational load. Within PTQ, we implement two approaches. In static quantization, model weights and activations are converted to lower precision prior to inference, whereas in dynamic quantization, only the weights are quantized while activations remain in floating-point representation.
Then, we incorporate Quantization-Aware Training (QAT). It is an advanced technique in which the effects of quantization are simulated during the training process. By integrating quantization constraints early in the learning pipeline, the model is trained to adapt to lower precision. This approach enhances the model’s robustness and generalization, resulting in a lower loss of accuracy compared to post-training quantization and enabling the model to better handle reduced precision for deployment across diverse hardware architectures.
Our implementation of quantization is carried out using the torch.quantization module in PyTorch. The quantization workflow involves three main stage. We utilize model preparation, quantization configuration, and calibration with evaluation. During model preparation, we select specific layers—such as linear projections and self-attention heads—that are critical for performance and amenable to quantization. In the quantization configuration step, high-precision floating-point values are replaced with integer-based representations (e.g., int8). Finally, we perform post-quantization calibration using validation data to ensure that any loss in accuracy is kept to a minimum.
@@
Computational Complexity and Energy Consumption
@@
Experimental Setup
System Configuration
@@
Implementation Details and Model Configurations
@@
Dataset Description
Table 1: Statistics of the Time Series Datasets
Dataset Train Size Test Size Length  Classes
PLAID 537 537 30 11
ElectricDevices 8926 7711 96 7
RefrigerationDevices 375 375 720 3
In our study, we utilize three datasets to analyze and classify electrical device usage patterns: RefrigerationDevices, ElectricDevices, and PLAID.
@@
Table 1 summarizes their key properties, including training/test sizes, sequence lengths, and number of classes.
Evaluation Metrics
In our experimental evaluation, we assessed the performance of our optimized transformer models using three primary metrics. First, we measured Classification Accuracy by calculating the proportion of correctly classified time series sequences, which served as a fundamental indicator of model performance. Second, we quantified Computational Overhead by evaluating reductions in inference time, the number of floating-point operations (FLOPs), and memory usage. These evaluations provided critical insights into the efficiency improvements achieved through our model optimization strategies. Finally, we determined Energy Efficiency by directly measuring the power consumption of each model during inference, thereby reflecting the effectiveness of our approaches in reducing energy consumption in practical deployments.
Results and Analysis
@@
Classification Accuracy Assessment
@@
Impact of Optimization Techniques
Through systematic application of our proposed optimization strategies, we observed that quantization techniques generally preserved model performance more effectively than pruning approaches, particularly in the context of the more complex T2 architecture. The implementation of 8-bit quantization resulted in a modest accuracy degradation of only 1.4-1.8% across all datasets, while achieving substantial improvements in computational efficiency and memory utilization, as detailed in subsequent sections of this analysis.
Energy Efficiency Evaluation
@@ 5.3.1 Computational Resource Utilization
The experimental results reveal significant improvements in computational efficiency through our optimization framework, with quantized models demonstrating reduced inference times ranging from 29.5% to 34.2% compared to their baseline counterparts. Furthermore, the integration of structured pruning techniques yielded additional performance benefits, particularly in memory-constrained environments, where we observed reductions in model size of up to 38.4% while maintaining classification accuracy within acceptable bounds of degradation.
Statistical Significance and Error Analysis
To ensure the statistical validity of our findings, we conducted comprehensive statistical analyses across multiple experimental runs, establishing confidence intervals at the 95% level through the application of the following statistical framework:
@@
Trade-off Analysis and Optimization Impact
The experimental results reveal meaningful insights regarding the trade-offs between model complexity, energy efficiency, and classification performance. Through careful analysis of these relationships, we observe that:
@@
Table 2: Comprehensive Performance Metrics Across Different Model Configurations and Datasets
Model Configuration Accuracy (%) Inference Time (ms) Energy (J) Memory (MB) FLOPs (G) RefrigerationDevices Dataset T1 Baseline 61.82 ± 0.45 6.42 ± 0.31 35.3 ± 2.1 689.5 4.82 T1 + Static Quantization 59.45 ± 0.52 4.35 ± 0.25 25.1 ± 1.8 172.4 4.82 T1 + Dynamic Quantization 58.28 ± 0.48 3.86 ± 0.28 23.5 ± 1.9 172.4 4.82 T1 + L1 Pruning 57.12 ± 0.55 3.41 ± 0.27 21.7 ± 1.7 275.8 2.89 T1 + L2 Pruning 56.45 ± 0.58 3.95 ± 0.29 22.8 ± 1.8 289.6 2.95 T2 Baseline 65.95 ± 0.38 7.32 ± 0.35 42.8 ± 2.3 1435.2 9.64 T2 + Static Quantization 63.24 ± 0.42 5.23 ± 0.28 31.4 ± 1.9 358.8 9.64 T2 + Dynamic Quantization 62.48 ± 0.45 4.96 ± 0.30 29.9 ± 2.0 358.8 9.64 T2 + L1 Pruning 61.72 ± 0.49 4.61 ± 0.31 27.5 ± 1.8 574.1 5.78 T2 + L2 Pruning 60.35 ± 0.51 5.03 ± 0.32 28.8 ± 1.9 602.8 5.89 ElectricDevices Dataset T1 Baseline 74.52 ± 0.42 11.01 ± 0.33 47.2 ± 2.2 689.5 4.82 T1 + Static Quantization 72.18 ± 0.48 7.89 ± 0.26 33.8 ± 1.9 172.4 4.82 T1 + Dynamic Quantization 71.35 ± 0.45 7.12 ± 0.29 31.9 ± 1.8 172.4 4.82 T1 + L1 Pruning 71.02 ± 0.52 6.79 ± 0.28 29.8 ± 1.8 275.8 2.89 T1 + L2 Pruning 70.45 ± 0.54 7.10 ± 0.30 30.9 ± 1.9 289.6 2.95 T2 Baseline 79.85 ± 0.35 12.78 ± 0.36 54.5 ± 2.4 1435.2 9.64 T2 + Static Quantization 78.12 ± 0.40 8.89 ± 0.29 39.8 ± 2.0 358.8 9.64 T2 + Dynamic Quantization 77.45 ± 0.43 8.12 ± 0.31 38.2 ± 1.9 358.8 9.64 T2 + L1 Pruning 76.82 ± 0.47 7.89 ± 0.32 35.9 ± 1.8 574.1 5.78 T2 + L2 Pruning 76.25 ± 0.49 8.30 ± 0.33 37.2 ± 2.0 602.8 5.89 PLAID Dataset T1 Baseline 80.95 ± 0.48 5.58 ± 0.32 36.1 ± 2.2 689.5 4.82 T1 + Static Quantization 78.42 ± 0.54 3.15 ± 0.27 24.8 ± 1.9 172.4 4.82 T1 + Dynamic Quantization 77.85 ± 0.51 2.75 ± 0.30 23.2 ± 1.8 172.4 4.82 T1 + L1 Pruning 77.52 ± 0.57 2.34 ± 0.29 21.1 ± 1.7 275.8 2.89 T1 + L2 Pruning 76.85 ± 0.59 2.85 ± 0.31 22.2 ± 1.8 289.6 2.95 T2 Baseline 84.25 ± 0.41 6.43 ± 0.37 43.2 ± 2.3 1435.2 9.64 T2 + Static Quantization 82.95 ± 0.45 4.45 ± 0.30 31.9 ± 2.0 358.8 9.64 T2 + Dynamic Quantization 82.12 ± 0.47 3.85 ± 0.32 30.1 ± 1.9 358.8 9.64 T2 + L1 Pruning 81.45 ± 0.52 3.45 ± 0.33 27.8 ± 1.8 574.1 5.78 T2 + L2 Pruning 80.92 ± 0.54 3.92 ± 0.34 29.1 ± 2.0 602.8 5.89
Table 3: Summary of Optimization Impacts
Optimization Method Accuracy Drop (%) Speed-up Energy Saving (%)
Static Quantization 2.37 ± 0.12 1.42× 29.14
Dynamic Quantization 3.14 ± 0.15 1.52× 33.25
L1 Pruning 3.43 ± 0.18 1.63× 37.08
L2 Pruning 3.82 ± 0.19 1.51× 35.12
Table 4: Energy-Accuracy Trade-off Analysis
Model Energy Efficiency Accuracy Retention Overall Score
(GFLOPS/J) (%) (EE × AR)
T1 Baseline 0.106 100.0 10.60
T1 + Static Quantization 0.150 96.8 14.52
T1 + L1 Pruning 0.168 95.2 15.99
T2 Baseline 0.182 100.0 18.20
T2 + Static Quantization 0.251 97.5 24.47
T2 + L1 Pruning 0.278 95.8 26.63
@@
Discussion
Impact on Classification Accuracy

### v4

@@ Abstract
The increasing computational demands of transformer models in time series classification necessitate effective optimization strategies for energy-efficient deployment. Our study presents a systematic investigation of optimization techniques, focusing on structured pruning and quantization methods for transformer architectures. Through extensive experimentation on three distinct datasets (RefrigerationDevices, ElectricDevices, and PLAID), model performance and energy efficiency are quantitatively evaluated across different transformer configurations. Our experimental results demonstrate that static quantization reduces energy consumption by 29.14% while maintaining classification performance, and L1 pruning achieves a 63% improvement in inference speed with minimal accuracy degradation. Our findings provide valuable insights into the effectiveness of optimization strategies for transformer-based time series classification, establishing a foundation for efficient model deployment in resource-constrained environments.
keywords:
Energy Efficiency, Time Series Classification, Optimization, Quantization, Pruning
1 Introduction
@@
Recent advancements in transformer architectures have significantly advanced time series analysis across various domains, including healthcare, finance, and predictive maintenance, enabling more accurate forecasting, anomaly detection, and decision-making processes [10, 13]. The ability of AI models to process sequential data efficiently has enabled substantial improvements in these fields, allowing for more accurate forecasting, anomaly detection, and decision-making processes [28, 23]. At the core of these developments are transformer-based architectures, which have demonstrated superior performance over traditional models, such as recurrent neural networks (RNNs) and long short-term memory (LSTM) networks, by leveraging self-attention mechanisms to capture long-range dependencies in temporal data [36]. Despite their effectiveness, transformers are computationally expensive, making them less viable for real-time and edge-based applications due to their high energy consumption and memory footprint. Furthermore, the growing carbon footprint associated with transformer training and inference has raised significant concerns regarding sustainability and deployment feasibility in resource-constrained environments [25, 31]. With the exponentially increasing usage of AI, the carbon footprint of massive models has been a topic of increasing worry. At the same time, deep learning model’s runaway scaling, such as huge transformers and diffusion models, has induced unmatched computation costs, which are enormous and require a lot of power to operate and train, directly contributing to the increase in global carbon emissions. Consequently, feeding environmental concerns such as global warming. If not checked, the long-term growth of AI systems will hugely frustrate efforts at sustainability and drain energy resources.
@@
focusing to encompass computation overheads, researchers are developing strategies not to compromise forecasting performance. Enhancing model efficiency, in addition to techniques like knowledge distillation, pruning, quantization, and even hardware-based architectures with increased energy efficiency, should be a pursuit by research professionals[7]. Furthermore, green AI technologies, i.e., renewable energy for data centers and minimizing the algorithms to consume less energy, need to be transitioned in order to make AI a force for good and not an environmental issue.
@@ 2 Literature Review
2.1 Transformers in Time Series Forecasting and Classification
@@
2.2 Model Compression Strategies
@@
2.3 Optimizing Transformer Inference
@@ 3.2 Data Preprocessing
3.3 Transformer Model Implementation
Our study incorporates the Vision Transformer (ViT) model [5], which has been increasingly adapted for time series tasks due to its ability to capture long-range dependencies efficiently [15]. Our model architecture comprises several key components. First, the Time Series Patch Embedding Layer transforms raw time series data into fixed-size patches using a one-dimensional convolutional layer, with positional encoding added to maintain the sequential order. Next, the multi-head self-attention mechanism allows the model to focus on different segments of the time series simultaneously, enhancing its ability to learn complex relationships. Each encoder block contains position-wise feed-forward layers with ReLU activations and dropout layers to reduce overfitting. To ensure stable training and mitigate the risk of vanishing gradients, the architecture incorporates layer normalization and residual connections. Finally, a fully connected softmax output layer classifies each sequence into its corresponding device category [29].
@@ 3.4.2 Quantization
where $Q$ denotes the quantization factor. Reducing numerical precision can lead to substantial power savings while maintaining a minimal impact on model performance. In our study, two distinct quantization strategies are employed to decrease memory footprint and computational overhead while maintaining high model accuracy.
First, Post-Training Quantization (PTQ) is applied. It reduces the precision of model weights and activations after training by converting 32-bit floating-point numbers into lower-bit representations. This process significantly decreases the model’s memory requirements and computational load. Within PTQ, we implement two approaches. In static quantization, model weights and activations are converted to lower precision prior to inference, whereas in dynamic quantization, only the weights are quantized while activations remain in floating-point representation.
Then, Quantization-Aware Training (QAT) is incorporated. It is an advanced technique in which the effects of quantization are simulated during the training process. By integrating quantization constraints early in the learning pipeline, the model is trained to adapt to lower precision. This approach enhances the model’s robustness and generalization, resulting in a lower loss of accuracy compared to post-training quantization and enabling the model to better handle reduced precision for deployment across diverse hardware architectures.
Our implementation of quantization is carried out using the torch.quantization module in PyTorch. The quantization workflow involves three main stage, model preparation, quantization configuration, and calibration with evaluation. During model preparation, specific layers—such as linear projections and self-attention heads—that are critical for performance and amenable to quantization are selected. In the quantization configuration step, high-precision floating-point values are replaced with integer-based representations (e.g., int8). Finally, post-quantization calibration is performed using validation data to ensure that any loss in accuracy is kept to a minimum.
@@
3.5 Computational Complexity and Energy Consumption
@@
4 Experimental Setup
4.1 System Configuration
@@
4.2 Implementation Details and Model Configurations
@@
4.3 Dataset Description
[…]
Properties of Time Series Datasets
Dataset Train Size Test Size Length  Classes PLAID 537 537 30 11 ElectricDevices 8926 7711 96 7 RefrigerationDevices 375 375 720 3
In this study, three datasets are utilized to analyze and classify electrical device usage patterns: RefrigerationDevices, ElectricDevices, and PLAID.
@@ 4.3 Dataset Description
Table LABEL:tab:datasets summarizes their key properties, including training/test sizes, sequence lengths, and number of classes.
4.4 Evaluation Metrics
In the experimental evaluation, the performance of our optimized transformer models is assessed using three primary metrics. First, Classification Accuracy is measured by calculating the proportion of correctly classified time series sequences, which serves as a fundamental indicator of model performance. Second, Computational Overhead is quantified by evaluating reductions in inference time, the number of floating-point operations (FLOPs), and memory usage. These evaluations provided critical insights into the efficiency improvements achieved through our model optimization strategies. Finally, Energy Efficiency is determined by directly measuring the power consumption of each model during inference, thereby reflecting the effectiveness of our approaches in reducing energy consumption in practical deployments.
5 Results and Analysis
@@
5.1 Classification Accuracy Assessment
@@
5.2 Impact of Optimization Techniques
Through systematic application of our proposed optimization strategies, it is observed that quantization techniques generally preserved model performance more effectively than pruning approaches, particularly in the context of the more complex T2 architecture. The implementation of 8-bit quantization resulted in a modest accuracy degradation of only 1.4-1.8% across all datasets, while achieving substantial improvements in computational efficiency and memory utilization, as detailed in subsequent sections of this analysis.
5.3 Energy Efficiency Evaluation
@@ 5.3.1 Computational Resource Utilization
The experimental results reveal significant improvements in computational efficiency through our optimization framework, with quantized models demonstrating reduced inference times ranging from 29.5% to 34.2%, compared to their baseline counterparts. Furthermore, the integration of structured pruning techniques yielded additional performance benefits, particularly in memory-constrained environments, where reductions in model size are observed of up to 38.4%, while maintaining classification accuracy within acceptable bounds of degradation.
5.4 Statistical Significance and Error Analysis
To ensure the statistical validity of the findings, comprehensive statistical analyses is conducted across multiple experimental runs, establishing confidence intervals at the 95% level through the application of the following statistical framework:
@@ 5.5 Comparative Analysis
5.6 Trade-off Analysis and Optimization Impact
The experimental results reveal meaningful insights regarding the trade-offs between model complexity, energy efficiency, and classification performance. Through careful analysis of these relationships, it is observe that:
@@ 5.6 Trade-off Analysis and Optimization Impact
[…]
Comprehensive Performance Metrics Across Different Model Configurations and Datasets Model Configuration Accuracy (%) Inference Time (ms) Energy (J) Memory (MB) FLOPs (G) RefrigerationDevices Dataset T1 Baseline 61.82 ± 0.45 6.42 ± 0.31 35.3 ± 2.1 689.5 4.82 T1 + Static Quantization 59.45 ± 0.52 4.35 ± 0.25 25.1 ± 1.8 172.4 4.82 T1 + Dynamic Quantization 58.28 ± 0.48 3.86 ± 0.28 23.5 ± 1.9 172.4 4.82 T1 + L1 Pruning 57.12 ± 0.55 3.41 ± 0.27 21.7 ± 1.7 275.8 2.89 T1 + L2 Pruning 56.45 ± 0.58 3.95 ± 0.29 22.8 ± 1.8 289.6 2.95 T2 Baseline 65.95 ± 0.38 7.32 ± 0.35 42.8 ± 2.3 1435.2 9.64 T2 + Static Quantization 63.24 ± 0.42 5.23 ± 0.28 31.4 ± 1.9 358.8 9.64 T2 + Dynamic Quantization 62.48 ± 0.45 4.96 ± 0.30 29.9 ± 2.0 358.8 9.64 T2 + L1 Pruning 61.72 ± 0.49 4.61 ± 0.31 27.5 ± 1.8 574.1 5.78 T2 + L2 Pruning 60.35 ± 0.51 5.03 ± 0.32 28.8 ± 1.9 602.8 5.89 ElectricDevices Dataset T1 Baseline 74.52 ± 0.42 11.01 ± 0.33 47.2 ± 2.2 689.5 4.82 T1 + Static Quantization 72.18 ± 0.48 7.89 ± 0.26 33.8 ± 1.9 172.4 4.82 T1 + Dynamic Quantization 71.35 ± 0.45 7.12 ± 0.29 31.9 ± 1.8 172.4 4.82 T1 + L1 Pruning 71.02 ± 0.52 6.79 ± 0.28 29.8 ± 1.8 275.8 2.89 T1 + L2 Pruning 70.45 ± 0.54 7.10 ± 0.30 30.9 ± 1.9 289.6 2.95 T2 Baseline 79.85 ± 0.35 12.78 ± 0.36 54.5 ± 2.4 1435.2 9.64 T2 + Static Quantization 78.12 ± 0.40 8.89 ± 0.29 39.8 ± 2.0 358.8 9.64 T2 + Dynamic Quantization 77.45 ± 0.43 8.12 ± 0.31 38.2 ± 1.9 358.8 9.64 T2 + L1 Pruning 76.82 ± 0.47 7.89 ± 0.32 35.9 ± 1.8 574.1 5.78 T2 + L2 Pruning 76.25 ± 0.49 8.30 ± 0.33 37.2 ± 2.0 602.8 5.89 PLAID Dataset T1 Baseline 80.95 ± 0.48 5.58 ± 0.32 36.1 ± 2.2 689.5 4.82 T1 + Static Quantization 78.42 ± 0.54 3.15 ± 0.27 24.8 ± 1.9 172.4 4.82 T1 + Dynamic Quantization 77.85 ± 0.51 2.75 ± 0.30 23.2 ± 1.8 172.4 4.82 T1 + L1 Pruning 77.52 ± 0.57 2.34 ± 0.29 21.1 ± 1.7 275.8 2.89 T1 + L2 Pruning 76.85 ± 0.59 2.85 ± 0.31 22.2 ± 1.8 289.6 2.95 T2 Baseline 84.25 ± 0.41 6.43 ± 0.37 43.2 ± 2.3 1435.2 9.64 T2 + Static Quantization 82.95 ± 0.45 4.45 ± 0.30 31.9 ± 2.0 358.8 9.64 T2 + Dynamic Quantization 82.12 ± 0.47 3.85 ± 0.32 30.1 ± 1.9 358.8 9.64 T2 + L1 Pruning 81.45 ± 0.52 3.45 ± 0.33 27.8 ± 1.8 574.1 5.78 T2 + L2 Pruning 80.92 ± 0.54 3.92 ± 0.34 29.1 ± 2.0 602.8 5.89
[…]
Summary of Optimization Impacts
Optimization Method Accuracy Drop (%) Speed-up Energy Saving (%) Static Quantization 2.37 ± 0.12 1.42× 29.14 Dynamic Quantization 3.14 ± 0.15 1.52× 33.25 L1 Pruning 3.43 ± 0.18 1.63× 37.08 L2 Pruning 3.82 ± 0.19 1.51× 35.12
[…]
Energy-Accuracy Trade-off Analysis
Model Energy Efficiency Accuracy Retention Overall Score (GFLOPS/J) (%) (EE × AR) T1 Baseline 0.106 100.0 10.60 T1 + Static Quantization 0.150 96.8 14.52 T1 + L1 Pruning 0.168 95.2 15.99 T2 Baseline 0.182 100.0 18.20 T2 + Static Quantization 0.251 97.5 24.47 T2 + L1 Pruning 0.278 95.8 26.63
@@
6 Discussion
6.1 Impact on Classification Accuracy